# Verb Mirage: Unveiling and Assessing Verb Concept Hallucinations in Multimodal Large Language Models

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have garnered significant attention recently and demonstrate outstanding capabilities in various tasks such as OCR, VQA, captioning, etc. However, hallucination remains a persistent issue. While numerous methods have been proposed to mitigate hallucinations, achieving notable improvements, these methods primarily focus on mitigating hallucinations about object/noun-related concepts. Verb concepts, crucial for understanding human actions, have been largely overlooked. In this paper, to the best of our knowledge, we are the first to investigate the verb hallucination phenomenon of MLLMs from various perspectives. Our findings reveal that most state-of-the-art MLLMs suffer from severe verb hallucination. To assess the effectiveness of existing mitigation methods for object concept hallucination on verb hallucination, we evaluated these methods and found that they do not effectively address verb hallucination. To address this issue, we propose a novel rich verb knowledge-based tuning method to mitigate verb hallucination. The experiment results demonstrate that our method significantly reduces hallucinations related to verbs. Our code and data will be made publicly available.
1 Introduction
@@
Multimodal Large Language Models (MLLMs) [zhu2024minigpt, dai2023instructblip, liu2024improved, chen2024internvl, bai2023qwen, ye2024mplug] have drawn much attention in both research and industry community. Armed with high-quality data, a large number of parameters, and efficient instruction-following finetuning [liu2023visual], they achieve great success in many tasks such as OCR, VQA, and image captioning, demonstrating strong generalization ability.
@@
Figure 1: Besides the well-discussed object hallucination, in this paper, we unveil the severe verb hallucination of state-of-the-art MLLMs with our designed benchmarks via multiple choice questions (MCQ) and binary questions (YN). Gemini-1.5-Flash and GPT-4-Turbo are tested with 100 randomly sampled questions.
@@
However, MLLMs’ performance improvement could be hindered by hallucination. Typically, hallucination [ji2023survey, liu2024survey] means the output of MLLM contains contents against facts, irrelevant or nonsensical given context such as prompt or multimodal input. To test MLLM hallucination in different tasks, many benchmarks [liu2024mmbench, guan2024hallusionbench, cai2023benchlmm, fu2023mme] have been made, allowing people to assess MLLMs’ abilities in various aspects. To mitigate MLLM hallucination, many methods [huang2024opera, leng2024mitigating, sun2023aligning, yin2023woodpecker, zhou2023analyzing] have been proposed, successfully relieving hallucination to a large extent.
@@
However, existing benchmarks and methods mainly target hallucination about objects/noun-related concepts. Action/verb-related concepts, which are crucial to understanding human actions, are typically overlooked. To this end, we propose to dig into the verb hallucination problem. We build the first verb-hallucination-oriented benchmark, which is based on existing datasets [chao2015hico, sigurdsson2018actor] needless of extra manual annotations. As MLLMs are a cooperation of vision and language modalities, we probe MLLM verb hallucination given both different visual inputs and language inputs, covering different query conditions, different imaging conditions, and different semantic conditions. Extensive experiments show that all MLLMs perform poorly on many aspects and thus show severe verb hallucination.
@@
    To our knowledge, we point out and analyze MLLM’s verb hallucination for the first time and probe this phenomenon from different aspects.
@@
    We evaluate existing low-cost hallucination mitigation methods on widely used MLLMs, and find out they all fail to alleviate verb hallucination.
@@
    We propose a baseline that can relieve verb hallucination better and discuss the reason for verb hallucination and possible solutions, paving the way for eliminating verb hallucination in future work.
2 Related Work
2.1 MLLM Benchmarks
@@
Before the emergence of LLMs, great efforts have been made to build datasets focused on specific tasks like image captioning [chen2015microsoft, sharma2018conceptual, young2014image], VQA [goyal2017making, hudson2019gqa, marino2019ok], OCR [singh2019towards], action recognition and detection [tincvpr, chao2015hico, djrn, idn, sigurdsson2018actor, gu2018ava, pastanet], etc. However, they mainly assess the abilities of domain-specific tasks. To fully evaluate the abilities of MLLMs, more benchmarks have been proposed [fu2023mme, liu2024mmbench, li2024seed, ying2024mmtbench] to test different aspects and subtasks. Benchmarks are also proposed to conduct detailed assessments on specific aspects, like BenchLMM [cai2023benchlmm] for robustness against image styles, MMSpuBench [ye2024mm] for robustness against spurious correlations, and HaloQuest [wang2024haloquest] for MLLM multimodal reasoning capabilities.
2.2 MLLM Hallucination
@@
Among the emerging benchmarks, hallucination has become a major focus. Typically, hallucination means that the contents generated by models are untruthful, against facts, or nonsensical [ji2023survey, liu2024survey, zhang2023siren]. Many benchmarks targeting hallucination have been proposed [guan2024hallusionbench, ye2024mm, wang2024haloquest]. In POPE [li2023evaluating], binary questions are used to probe hallucination about a certain class of object. CHAIR score [rohrbach2018object] is used to measure object hallucination of MLLMs in image captioning. Researchers also study other phenomena, such as event hallucination [jiang2024hal], hallucination snowballing [zhong2024investigating], and relation hallucination [wu2024evaluating]. The cause of hallucination is also investigated. OPERA [huang2024opera] recognizes MLLMs’ over-attention to summarizing text tokens as an outstanding phenomenon of object hallucination. VCD [leng2024mitigating] reveals that language prior is an important factor in hallucination. However, in most previous methods, only hallucinations involving objects were covered by identifying whether the output of MLLM contains a reference to an object that does not match the image or appears in the image. Verb-related concepts, which are crucial to understanding human activities, were neglected. Instead, we omit relevance to object hallucination, focus on verb concepts, and probe verb hallucination on different vision and language conditions.
2.3 Hallucination Mitigation
@@
Researchers have revealed the reason for hallucination from many different aspects and proposed hallucination mitigation methods respectively. To omit the bias or errors in training data, researchers proposed mitigating bias [liu2024mitigating, hu2023ciem, gunjal2024detecting] in the dataset or enriching the annotation [zhai2023halle]. Some works mitigated hallucination by scaling up the resolution of input images [bai2023qwen, chen2024internvl, li2023monkey]. Moreover, some suggest post-processing at inference time by adjusting decoding strategy [huang2024opera, leng2024mitigating, chen2424halc] or correcting the output of MLLMs with the help of expert models [zhou2023analyzing, yin2023woodpecker].
3 Probing on MLLM IO Conditions
@@
We probe verb hallucination from different perspectives, e.g., MLLM behavior given different question formats, image qualities, verb semantics, and angles of view, etc.
@@
We select HICO [chao2015hico] and CharadesEgo [sigurdsson2018actor] as the main datasets for probing MLLM verb hallucination. HICO contains 47K images with dense annotations. It includes 600 action classes formed by 80 object classes and 117 verb classes. It has rich verb labels and is thus suitable for evaluating MLLM verb understanding with minor manual adaptation. Meanwhile, CharadesEgo contains 7K videos of daily indoor activities. In each scenario, the same actor is recorded with both an egocentric and exocentric camera. It contains 157 action classes, each formed by a verb and an object.
@@
We test several open-sourced and close-sourced MLLMs including InstructBLIP-7B [liu2023visual], LLaVA-V1.5-7B [liu2024improved], mPLUG-Owl2 [ye2024mplug], Qwen-VL-Chat [bai2023qwen], MiniCPM-Llama3-V2.5 [yao2024minicpm], Qwen2-VL-7B-Instruct [Qwen2VL], GPT-4-Turbo [achiam2023gpt], Gemini-1.5-Flash [team2023gemini], etc. They have different ranks on commonly used leaderboards and show outstanding results on benchmarks targeting object concepts.
@@
Figure 2: We probe MLLM verb hallucination from various perspectives, e.g., question formats, the existence of object correlation, different fields of view, image qualities, verb semantics, and image semantics.
3.1 Probing on Different Query Conditions
3.1.1 Different Question Formats
@@
Nearly all MLLMs follow the VQA format. Therefore, users can interact with MLLMs naturally by asking questions and requiring answers from MLLMs in natural language. MLLMs with low hallucinations should give hallucination-free answers given different question formats. Therefore, we evaluate MLLM hallucination using different question formats including Multiple Choice (MC) questions with only one correct answer each and Yes-or-No (YN) questions. Here we do not introduce free-form image captioning and blank-filling because these two forms require rule-based post-processing on verbs and may lead to severe misclassification errors.
When building MC questions, for a sample image, we randomly choose a verb presented in the image and three verbs possibly performable upon objects but not presented. We introduce circular evaluation [liu2024mmbench] and regard accuracy as a metric for MLLM verb hallucination. For YN questions, when building negative samples, we also randomly choose possible verbs for the objects but not carried out in the image. We regard accuracy, precision, recall, and F1 score as the metrics for hallucination.
@@
We aim to verify none other than verb hallucination, so we do our best to omit relevance to object hallucination. For each verb-object tuple, we form questions by altering the verb and leaving the object unchanged. For example, if an image contains a person holding a cup, we may ask MLLM, “Is someone holding a cup? Is someone washing a cup?” In this way, relevance with object hallucination can be minimized. However, we must point out that as a substantial proportion of verbs are transitive verbs, the influence of objects can not be completely omitted.
3.1.2 Object Correlation
@@
Sometimes we focus on human interaction with a certain class of object, but sometimes our focus on verbs may be object class agnostic. Specifically, we may wonder “Is someone holding a cup in the image?” However, sometimes we may also want to know “Is someone eating something in the image?” Therefore, we test MLLM verb understanding both given reference to objects and not. Among these two conditions, we believe questions without object correlation (i.e., “Is someone eating something?”) have less relevance to object hallucination, but questions with object correlation are also very practicable in daily use.
3.1.3 Analysis and Discussion
@@
Heavy reliance on objects. MLLMs show drastic performance degradation on MC questions without reference to objects. Detailed statistics on YN questions based on object classes referred to in the questions also reveal that MLLM verb understanding relies heavily on object reference. We analyze some commonly used datasets [sharma2018conceptual, changpinyo2021conceptual, lin2014microsoft, ordonez2011im2text, krishna2017visual, hudson2019gqa, goyal2017making] for MLLM pretraining and investigate the number of nouns and verbs in the datasets in Fig. 3(a). We can see that the number of nouns is 4-10 times the number of verbs. One reason for this unbalanced ratio of nouns and verbs is that datasets on action understanding have not attracted enough attention from the MLLM community. Another reason is that there are many more nouns than verbs in English. Research on shortcut learning [geirhos2020shortcut] also sheds light on the reason for MLLMs’ overreliance on nouns.
@@
3.2 Probing on Different Imaging Conditions
3.2.1 High-Quality and Low-Quality Images
Previous research [li2022hake, leng2024mitigating] has revealed that visually distorted images can hinder both humans and models from recognizing the contents in images well. However, the relation between visual distortion and verb hallucination is unexplored. Do MLLMs hallucinate in the same way when given high-quality images and visually distorted images? Is verb understanding more sensitive to visual distortion than object understanding for MLLMs? Here, we add pepper-salt noise as visual distortion to images for testing.
We evaluate MLLM verb understanding with both high-quality and visually-distorted images and report performance and error consistency following [geirhos2021partial] in Tab. 2. All tested MLLMs show obvious performance degradation. Error consistency in the form of Cohen’s Kappa [cohen1960coefficient] measures MLLM consistency of answers given different visual conditions and provides a guideline for MLLM performance improvement. We can see some MLLMs with low ranks do not have bad error consistency, but MLLMs with higher ranks show low error consistency. This means that their verb hallucination can be easily induced by visual distortion.
As a control experiment, besides verb understanding, we also build a test set for MLLMs’ object understanding with the same set of images. The observation consistency in terms of Cohen’s Kappa is reported in Fig. 3(b). From the result, we can see that MLLM shows much higher inconsistency in verb understanding than object understanding, which means that visual distortion does more harm to verb understanding than object understanding.
In conclusion, we have the following findings:
Finding: Visual distortion affects both object understanding and verb understanding of MLLMs, but the effect on verb understanding is greater.
@@ 3.2.1 High-Quality and Low-Quality Images
YN acc YN prec YN recall YN acc YN prec YN recall MC acc MC acc
Qwen2-VL-7B 74.69 57.43 95.87 63.20 47.68 95.57 56.86 65.31 51.94 48.21
MiniCPM-Llama3-V2.5 79.14 63.33 90.33 67.40 51.25 64.79 26.12 60.77 40.50 37.20
Qwen-VL-Chat 79.24 65.09 82.68 66.64 50.28 80.43 38.47 54.57 33.98 43.38
LLaVA V1.5 58.21 44.49 97.49 51.29 40.66 97.30 73.85 51.00 49.97 68.37
InstructBLIP 73.82 57.25 87.79 71.04 54.35 87.40 74.16 6.25 6.34 82.33
@@
3.2.2 Egocentic and Exocentric Images
@@
Given the same scenario recorded with exocentric and egocentric cameras, we also probe MLLMs’ understanding with MC and YN questions. The results are shown in Fig. 4 and Tab. 3, there is a substantial gap between MLLMs’ understanding of exocentric and egocentric images.
Finding: MLLMs can not understand verb concepts in egocentric images as well as exocentric images. MLLMs are better at exocentric verb understanding.
@@
3.3 Probing on Different Semantic Conditions
3.3.1 Rare and Common Verbs
Verbs follow long-tailed distribution in action datasets because of all sorts of difficulties in the process of dataset collection. However, understanding rare and common verbs is equally important in real-world applications. We hypothesize that MLLMs tend to hallucinate more on rare verbs and try to prove it on existing datasets.
We divide the negative samples into two subsets: the rare set and the common set. In rare set, the verb in question lies in the tail of HICO verb distribution while in common set, the verb in question lies in the head. Specifically, for YN questions with an object reference, the rare set contains all HOIs whose annotations make up less than 20% among HOIs relevant to the same object class. For YN questions without object reference, the common set contains all questions containing hold, ride, sit_on, straddle and carry, making up 50% of the verb annotations in HICO dataset. From Fig. 5, we can see that MLLMs tend to refuse existent rare verbs but accept nonexistent common verbs in images. This phenomenon reveals that the long-tailed distribution of verb annotations limits MLLM verb understanding. How to understand rare verbs remains a problem and there is a large room for action data collection and curation.
Finding: MLLMs can not understand rare verbs as well as common verbs, i.e., long-tail affects a lot.
Figure 5: True negative correct rate comparison on rare and common subsets.
3.3.2 Image Content Ambiguity
@@
  • Imbalanced human-object relative size. Imbalanced human-object relative sizes can add difficulties to MLLM perception of humans and objects. The existence of verbs relies heavily on the accurate perception of humans and objects in images. Potential failures in perception bring great difficulties to the recognition of verbs.
  • Crowded scene. A highly complicated scene structure can distract MLLMs. To judge the existence of a verb, MLLMs must analyze all humans and objects and get a comprehensive conclusion. The large number of humans and objects puts heavy burdens on MLLMs.
  • Occlusion. We can recognize humans and objects according to part of them and analyze their relationship even with prominent occlusion. Thus, visual reasoning under occluded scenarios is important in fields like action understanding and robot manipulation.
From Fig. 6, we have the following finding:
Finding: MLLMs show performance degradation when images contain content ambiguity. How to understand verbs given ambiguous content is an open problem to be solved.
@@
4 Probing on Model Behaviors
In this section, we probe verb hallucination from the model perspective.
4.1 Token Uncertainty
We first dig into the uncertainty [zhou2023analyzing] of MLLMs via visualizing the distribution of probabilities of predicted tokens of the widely used open-sourced LLaVA V1.5. From Fig. 7, we can see that there is a substantial difference between correct answers and hallucinated answers: hallucinated tokens are mostly given with low probability, which means that the model is confused about the answers. For Yes-or-No questions, we visualize questions to which the model gives answers Yes and No separately. The observations are:

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have garnered significant attention recently and demonstrate outstanding capabilities in various tasks such as OCR, VQA, captioning, etc. However, hallucination remains a persistent issue. While numerous methods have been proposed to mitigate hallucinations, achieving notable improvements, these methods primarily focus on mitigating hallucinations related to object/noun concepts. Verb concepts, which are crucial for understanding human actions, have been largely overlooked. In this paper, to the best of our knowledge, we are the first to investigate the verb hallucination phenomenon of MLLMs from various perspectives. Our findings reveal that most state-of-the-art MLLMs suffer from severe verb hallucination. To assess the effectiveness of existing mitigation methods for object concept hallucination in relation to verb hallucination, we evaluated these methods and found that they do not effectively address verb hallucination. To address this issue, we propose a baseline method based on fine-tuning with rich verb knowledge, achieving decent superiority. The experiment results demonstrate that our method significantly reduces hallucinations related to verbs.
Codehttps://github.com/davidwang200099/Verb˙Mirage
Introduction
@@
Multimodal Large Language Models (MLLMs) (zhu2024minigpt; liu2023visual; chen2024internvl; bai2023qwen; li2024monkey) have drawn much attention in both the research and industry communities. Armed with high-quality data, a large number of parameters, and efficient instruction-following fine-tuning, they achieve significant success in various tasks, including OCR, VQA, and image captioning, demonstrating a strong generalization ability.
@@
Figure 1: Besides the well-discussed object hallucination, in this paper, we unveil the severe verb hallucination of state-of-the-art MLLMs with our designed benchmarks. All models show low object hallucination (on POPE) but severe verb hallucination. Gemini-1.5-Flash and GPT-4-Turbo are tested with 100 randomly sampled questions.
@@
However, MLLMs’ performance improvement could be hindered by hallucination. Typically, hallucination (ji2023survey; liu2024survey) means the output of MLLM contains content against facts, irrelevant or nonsensical given context, such as prompt or multimodal input. To test MLLM hallucination in different tasks, many benchmarks (liu2024mmbench; cai2023benchlmm; fu2023mme) have been made, allowing people to assess MLLMs’ abilities in various aspects. To mitigate MLLM hallucination, many methods (huang2024opera; leng2024mitigating; sun2023aligning; yin2023woodpecker; zhou2023analyzing) have been proposed, successfully relieving hallucination to a large extent.
However, existing benchmarks and methods mainly target hallucination about objects/noun-related concepts. Action/verb-related concepts, which are crucial to understanding human events, are overlooked.
@@
To this end, we propose to dig into the verb hallucination problem. We build the first verb-hallucination-oriented benchmark, which is based on existing datasets (chao2015hico; sigurdsson2018actor) without the need for extra manual annotations. As MLLMs are a cooperation of vision and language modalities, we probe MLLM verb hallucination given both different visual inputs and language inputs, covering different query conditions, different imaging conditions, and different semantic conditions. Extensive experiments show that all MLLMs perform poorly on many aspects and thus show severe verb hallucination.
@@
    To our knowledge, we point out and analyze MLLM’s verb hallucination for the first time and probe this phenomenon under different conditions.
@@
    We study the influence of existing training-free and finetuning-based methods on MLLM verb hallucination. We find that fine-tuning is still the most promising way for verb hallucination mitigation.
@@
    We probe model behavior from the perspective of vision-language interaction and token uncertainty, and study how well MLLMs learn verbs.
Related Work
@@
MLLM Benchmarks. Before the emergence of MLLMs, great efforts have been made to build datasets on tasks like image captioning (chen2015microsoft; sharma2018conceptual; young2014image), VQA (goyal2017making; hudson2019gqa; marino2019ok), OCR (singh2019towards), action recognition and detection (gu2018ava; tincvpr; djrn; idn; pastanet), etc. However, they mainly assess domain-specific models. To fully evaluate MLLMs, more benchmarks have been proposed (liu2024mmbench; ying2024mmtbench; yue2024mmmu; li2024seed; tong2024eyes; li2025the) to test different aspects and subtasks. Benchmarks are also proposed to conduct detailed assessments on specific aspects, like BenchLMM (cai2023benchlmm) for robustness against image styles and MMSpuBench (ye2024mm) for robustness against spurious correlations.
@@
MLLM Hallucination. Among the emerging benchmarks, hallucination has become a focus. Typically, hallucination means that the contents generated by models are untruthful, against facts, or nonsensical (ji2023survey; liu2024survey; zhang2023siren). Many benchmarks on hallucination have been proposed (guan2024hallusionbench; kaul2024throne; wang2024haloquest; chen2024multi; zhong2024investigating). Among different types of hallucination, object hallucination (rohrbach2018object) is deeper explored. Binary questions are used to probe hallucination about a certain class of objects in POPE (li2023evaluating). CHAIR score (rohrbach2018object) is used to measure object hallucination in image captioning. In most previous studies, only object concept hallucinations were covered by identifying whether MLLMs refer to objects nonexisting in the image or incorrect attributes of objects. Though some other phenomena are studied, such as event hallucination (jiang2024hal), relationship hallucination (wu2024evaluating), etc., verb-related concepts, which are crucial to understanding human actions, were still neglected. Event is a complex combination of objects, verbs, adjectives, adverbs, etc., and Hal-eval (jiang2024hal) tests event hallucination by introducing nonexisting objects and can be bypassed by mitigating object hallucination. Verbs are a kind of relation, but visual relationship also covers spatial relations, ownership, subject-place relations, attributes, etc., which account for the majority. In this paper, we focus on the understanding of humans existing in the image, i.e., a human-centric problem.
@@
Hallucination Mitigation. Researchers have revealed the reasons for hallucination from many different aspects and proposed hallucination mitigation methods. Over-attention to summarizing tokens (huang2024opera) and language prior (leng2024mitigating) are recognized to be correlated with object hallucination. To mitigate the bias or errors in training data, researchers proposed mitigating bias (liu2024mitigating; hu2023ciem; gunjal2024detecting) in the dataset or enriching the annotation (zhai2023halle). Moreover, some suggest post-processing at inference time by adjusting decoding strategy (leng2024mitigating; chen2424halc) or correcting the output of MLLMs with the help of expert models (zhou2023analyzing; yin2023woodpecker).
Probing on MLLM IO Conditions
@@
As shown in Fig. 2, we probe verb hallucination from various perspectives: MLLM behavior given different question formats, image qualities, verb semantics, viewpoints, etc.
@@
We select HICO (chao2015hico) and CharadesEgo (sigurdsson2018actor) as the main datasets for probing verb hallucination. HICO contains 47K images with dense annotations. It includes 600 action classes formed by 80 object classes and 117 verb classes. WordNet (miller1995wordnet) was used to handle synonym and hierarchy problem. It has rich verb labels and is thus suitable for evaluating MLLM verb understanding with minor manual adaptation. Meanwhile, CharadesEgo contains 7K videos of daily indoor activities. In each scenario, the same actor is recorded with both an egocentric and an exocentric camera. It contains 157 action classes, each formed by a verb and an object.
@@
We test several open-sourced and close-sourced MLLMs including InstructBLIP-7B (dai2023instructblip), LLaVA-V1.5-7B (liu2024improved), mPLUG-Owl2 (ye2024mplug), Qwen-VL-Chat (bai2023qwen), MiniCPM-Llama3-V2.5 (yao2024minicpm), Qwen2-VL-7B-Instruct (Qwen2VL), Molmo-7B-D (deitke2025molmo), GPT-4 (achiam2023gpt), Gemini-1.5 (team2023gemini), etc. They have different ranks on leaderboards and show outstanding results on benchmarks targeting object concepts.
@@
Figure 2: We probe MLLM verb hallucination from various perspectives, eg., question formats, the existence of object correlation, different fields of view, image qualities, verb semantics, and image semantics.
Probing on Query Conditions
Question Formats
@@
Next, we probe the relation between verb hallucination and QA format. MLLMs with low hallucinations should give hallucination-free answers, given different question formats. Thus, we evaluate verb hallucination using different question formats, including Multiple Choice (MC) questions with only one correct answer each and Yes-or-No (YN) questions. Here, we do not introduce free-form image captioning and blank-filling because these two forms require rule-based post-processing on verbs and may lead to severe misclassification errors.
@@
We aim to verify none other than verb hallucination, so we do our best to omit relevance to object hallucination. For each verb-object tuple in the labels, we form questions or options by altering the verb and leaving the object unchanged. For YN questions, when building negative samples, we randomly choose verbs that are plausible for the objects but not carried out in the image. For example, if an image contains a person holding a cup, we may ask MLLM, “Is someone holding a cup? Is someone washing a cup?” We regard accuracy, precision, recall, and F1 score as the metrics for hallucination. Similarly, when building MC questions, for a sample image, we randomly choose a verb presented in the image and three verbs possibly performable upon objects but not presented. We introduce circular evaluation (liu2024mmbench) and regard accuracy as a metric for MLLM verb hallucination. Then, the relevance of object hallucination can be minimized. However, we must point out that as a substantial proportion of verbs are transitive verbs, the influence of objects can not be completely omitted. The construction of benchmarks is detailed in Suppl. Sec. 1.
Object Correlation
@@
Sometimes we focus on human interaction with a certain class of object, but sometimes our focus on verbs may be object class agnostic. Specifically, we may wonder “Is someone holding a cup in the image?” However, sometimes we may also want to know “Is someone eating something in the image?” Therefore, we test MLLM verb understanding, given reference to objects and not. Among these two conditions, we believe questions without object correlation (ie, “Is someone eating something?”) have less relevance to object hallucination. Still, questions with object correlation are also very practical in daily use.
Analysis
@@
Heavy reliance on objects. MLLMs show drastic performance degradation on MC questions without reference to objects. Detailed statistics on YN questions based on object classes referred to in the questions also reveal that MLLM verb understanding relies heavily on object reference. We analyze some commonly used datasets (sharma2018conceptual; changpinyo2021conceptual; lin2014microsoft; ordonez2011im2text; krishna2017visual; hudson2019gqa; goyal2017making) for MLLM pretraining and investigate the number of nouns and verbs in the datasets in Fig. 4(a). We can see that the number of nouns is 4-10 times the number of verbs. One reason for this unbalanced ratio of nouns and verbs is that datasets on action understanding have not attracted enough attention from the MLLM community. Another reason is that there are many more nouns than verbs in English. Research on shortcut learning (geirhos2020shortcut) also sheds light on the reason for MLLMs’ overreliance on nouns.
@@
Similar Bias. We ensemble the answers given by three outstanding models and show results in Tab. 3. There is no substantial improvement over the individual models, showing that the models share similar biases.
@@
Figure 3: Comparison of YN questions with correct answer No on rare and common subsets.
@@
Probing on Imaging Conditions
High-Quality and Low-Quality Images
Previous research (li2022hake; leng2024mitigating) has revealed that visually distorted images can hinder both humans and models from recognizing the contents in images well. However, the relation between visual distortion and verb hallucination is unexplored. Do MLLMs hallucinate in the same way when given high-quality images and visually distorted images? Is verb understanding more sensitive to visual distortion than object understanding for MLLMs? Here, we add pepper-salt noise as a visual distortion to images, each pixel of which is affected at a probability of 75%.
We evaluate MLLM verb understanding with both high-quality and visually-distorted images and report performance and error consistency following (geirhos2021partial) in Tab. 2. All tested MLLMs show obvious performance degradation. Error consistency in the form of Cohen’s Kappa (cohen1960coefficient) measures MLLM consistency of answers given different visual conditions and provides a guideline for MLLM performance improvement. We can see that some MLLMs with low ranks do not have bad error consistency, but MLLMs with higher ranks show low error consistency. This means that their verb hallucination can be easily induced by visual distortion.
As a control experiment, besides verb understanding, we also built a test set for MLLMs’ object understanding with the same set of images. The observation consistency in terms of Cohen’s Kappa is reported in Fig. 4(b). From the result, we can see that MLLM shows much higher inconsistency in verb understanding than object understanding, which means that visual distortion does more harm to verb understanding than object understanding. We attribute this result to the sparsity of verb semantics in pixel space.
In conclusion, visual distortion affects both object understanding and verb understanding of MLLMs, but the effect on verb understanding is greater.
@@
YN acc YN F1 YN acc YN F1 MC acc MC acc
MiniCPM-Llama3-V2.5 79.14 74.46 67.40 57.23 26.12 60.77 40.50 37.20
Qwen-VL-Chat 79.24 72.84 66.64 61.88 38.47 54.57 33.98 43.38
LLaVA V1.5 59.16 61.79 51.29 57.35 73.85 51.00 49.97 68.37
InstructBLIP 73.82 69.30 71.04 67.02 74.16 6.25 6.34 82.33
@@
YN w/ obj YN verb only
acc prec recall acc prec recall
MiniCPM 80.91 66.83 85.41 79.14 63.33 90.33
Qwen-VL-Chat 78.06 62.37 87.02 79.24 65.09 82.68
InstructBLIP 72.53 55.79 86.77 73.83 57.26 87.80
Ensemble 75.59 58.74 91.17 79.81 64.48 89.12
Table 3: Ensembled accuracy/precision/recall on YN question w/wo object reference.
Egocentric and Exocentric Images
@@
Given the same scenario recorded with exocentric and egocentric cameras, we also probe MLLMs’ understanding with MC and YN questions. The results are shown in Fig. 5 and Tab. 4. There is a substantial gap between MLLMs’ understanding of exocentric and egocentric images.
In conclusion, MLLMs can not understand verb concepts in egocentric images as well as exocentric images. MLLMs are better at exocentric verb understanding.
@@
Probing on Semantic Conditions
Rare and Common Verbs
Verbs follow a long-tailed distribution in action datasets because of all sorts of difficulties in the process of dataset collection. However, understanding rare and common verbs is equally important in real-world applications. We hypothesize that MLLMs tend to hallucinate more on rare verbs and try to prove it on existing datasets.
We divide the negative samples into two subsets: the rare set and the common set. In the rare set, the verb in question lies in the tail of HICO verb distribution, while in the common set, the verb in question lies in the head. Specifically, for YN questions with an object reference, the rare set contains all HOIs whose annotations make up less than 20% among HOIs relevant to the same object class. For YN questions without object reference, the common set contains all questions containing hold, ride, sit_on, straddle, and carry, making up 50% of the verb annotations in the HICO dataset. From Fig. 3, we can see that MLLMs tend to refuse existent rare verbs but accept nonexistent common verbs in images. This phenomenon reveals that the long-tailed distribution of verb annotations limits MLLM verb understanding. How to understand rare verbs remains a problem, and there is a large room for action, data collection, and curation. In conclusion, MLLMs can not understand rare verbs as well as common verbs, i.e., long-tail affects a lot.
Image Content Ambiguity
@@
  • Imbalanced human-object relative size. Imbalanced human-object relative sizes can add difficulties to MLLM’s perception of humans and objects. The existence of verbs relies heavily on the accurate perception of humans and objects in images. Potential failures in perception bring great difficulties to the recognition of verbs.
  • Crowded scene. A highly complicated scene structure can distract MLLMs. To judge the existence of a verb, MLLMs must analyze all humans and objects and draw a comprehensive conclusion. The large number of humans and objects puts heavy burdens on MLLMs.
  • Occlusion. We can recognize humans and objects according to parts of them and analyze their relationship, even with prominent occlusion. Thus, visual reasoning under occluded scenarios is important in fields like action understanding and robot manipulation.
From Fig. 6, we have the following finding: MLLMs show performance degradation when images contain content ambiguity. How to understand verbs given ambiguous content is an open problem to be solved.
@@
Probing on Model Behaviors
In this section, we take LLaVA V1.5 as an example and study the relationship between its behavior and verb hallucination. Our study finds that although verb hallucination shares some commonality with object hallucination, they are fundamentally different.
Vision-Language Interaction
Key Image Area Attention. First, we hypothesize that models pay less attention to key information in images and text when they give hallucinated answers. From Fig. 7(a), we can see that there is an obvious distinction of distribution between hallucinated attention and non-hallucinated attention, showing a strong correlation between inadequate attention to key areas and hallucination.
Visual Token Attention. The Visual Token Attention is defined as
where $\alpha_{ij}$ represents the attention weight assigned to token $i$ at head $j$ in the last transformer layer, $V$ denotes the set of visual tokens, and $T$ denotes the set of textual tokens. We visualize it in Fig. 8 and we find:
  1. 1.
    For questions with correct answer No hallucinated models tend to give more attention to visual tokens.
  2. 2.
    For MC questions, there is no obvious difference in vision token attention between hallucinated models and non-hallucinated models.
Therefore, unlike what has been found on object hallucination, more attention to visual tokens does not exclude verb hallucination.
Figure 7: Attention for YN questions with object references.
Figure 8: Probability distribution of visual token attention.
Token Uncertainty
Token Uncertainty and Hallucination
We dig into the uncertainty (zhou2023analyzing) of MLLMs via visualizing the distribution of probabilities of predicted tokens of the widely-used open-sourced LLaVA V1.5. From Fig. 9, we can see that there is a substantial difference between correct answers and hallucinated answers: hallucinated tokens are mostly given with low probability, which means that the model is confused about the answers. For Yes-or-No questions, we visualize questions to which the model gives answers Yes and No separately. The observations are: